### v2

# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

Abstract
Objective: To improve prediction of Chronic Kidney Disease (CKD) progression to End Stage Renal Disease (ESRD) using machine learning (ML) and deep learning (DL) models applied to an integrated clinical and claims dataset of varying observation windows, supported by explainable AI (XAI) to enhance interpretability and reduce bias.
Materials and Methods: We utilized data about 10,326 CKD patients, combining their clinical and claims information from 2009 to 2018. Following data preprocessing, cohort identification, and feature engineering, we evaluated multiple statistical, ML and DL models using data extracted from five distinct observation windows. Feature importance and Shapley value analysis were employed to understand key predictors. Models were tested for robustness, clinical relevance, misclassification errors and bias issues.
Results: Integrated data models outperformed those using single data sources, with the Long Short-Term Memory (LSTM) model achieving the highest AUC (0.93) and F1 score (0.65). A 24-month observation window was identified as optimal for balancing early detection and prediction accuracy. The 2021 eGFR equation improved prediction accuracy and reduced racial bias, notably for African American patients.
Discussion: Improved ESRD prediction accuracy, results interpretability and bias mitigation strategies presented in this study have the potential to significantly enhance CKD and ESRD management, support targeted early interventions and reduce healthcare disparities.
Conclusion: This study presents a robust framework for predicting ESRD outcomes in CKD patients, improving clinical decision-making and patient care through multi-sourced, integrated data and AI/ML methods. Future research will expand data integration and explore the application of this framework to other chronic diseases.
1 Introduction
[…]
Previous efforts to predict CKD progression to ESRD have primarily relied on either clinical data from electronic health records (EHRs) 4 or administrative claims data 5. While informative, these studies often utilize a limited set of features, potentially under-representing the complexity of CKD progression. For instance, Sharma et al. 6 developed a model using claims data to identify CKD patients at risk of Hyperkalemia, and Krishnamurthy et al. 5 predicted CKD onset using comorbidities and medications from Taiwan’s National Health Insurance Database. However, claims-based studies generally lack the granularity of clinical data, which offers detailed insights through laboratory results, medication records, and patient demographics. For example, Tangri et al. 7 used clinical variables like age, gender, eGFR, and albuminuria to predict ESRD progression, and Keane et al. 8 utilized serum creatinine and urine protein levels for high-risk identification. Despite their value, models based solely on clinical data may miss the full spectrum of patient interactions with the healthcare system and often struggle with missing data and inconsistent recording practices. Additionally, they may overlook socio-economic factors and healthcare utilization patterns, limiting their applicability across diverse patient populations.
In the field of explainable AI (XAI) for healthcare, established methods like SHapley Additive exPlanations (SHAP) 9 and Local Interpretable Model-agnostic Explanations (LIME) 10 have been applied to predict ESRD, laying a strong foundation for enhancing the interpretability of deep learning models. However, relying solely on a single source of data — whether claims or clinical — restricts the ability to fully understand the complexity and variability of CKD progression across a diverse patient population and predict ESRD, thereby limiting the real-world applicability of these models.
This study aims to bridge this gap by developing a robust and comprehensive framework that not only utilizes clinical or claims data in isolation but also harnesses the combined strengths of these integrated data sets. At the same time, by minimizing the observation window needed for accurate prediction, our framework integrates clinical relevance with a practical, patient-centered focus. This approach enhances predictive accuracy and clinical utility, enabling more informed decision-making and ultimately improving patient outcomes.
2 Objective
This research aims to evaluate multiple prediction models for CKD progression to ESRD using an integrated dataset of clinical and claims data. Our primary objective is to estimate the probability that a CKD patient will develop ESRD in the future based on the data from a specified observation window prior to the diagnosis of ESRD, represented as follows:
$P\left(\mathrm{ESRD}_{i}=1\mid t,X_{\text{clinical},i}(t),X_{\text{claim},i}(t)\right),$
where $X_{\text{clinical},i}(t)$ and $X_{\text{claim},i}(t)$ denote the clinical and claims data observed up to time $t$ for patient $i$ who progresses to ESRD after the observation window.
We address two key scenarios in our study:
  • Model Robustness: We train and test different analytical models on the same dataset to ensure consistency and comparability, allowing us to select the best-performing model with a focus on robustness and reliability.
  • Clinical Utility: In clinical settings, the likely duration of CKD progression to ESRD is unknown. Our model aims to predict the likelihood of ESRD in the future and suggests an optimal observation window based on the duration of data available on the patient from the time of CKD Stage 3 diagnosis in time steps of 6, 12, 18, 24 and 30 months.
To address these objectives, we focus on identifying the best data-driven observation window while maintaining an acceptable error rate $\epsilon(t)$, represented as follows:
$T_{\text{obs }}=\min\left\{t\mid\epsilon(t)\leq\epsilon^{\prime},t\in\{6,12,18,24,30\}\right\},\vspace{-.5cm}$
where $\epsilon^{\prime}$ is defined as:
$\epsilon^{\prime}=\left\{\begin{array}[]{ll}\textnormal{user-defined error rate}&\textnormal{if specified},\\ \min_{t\in\{6,12,18,24,30\}}\epsilon(t)&\textnormal{otherwise}\end{array}\right.$
Additionally, we apply explainable AI techniques, such as feature importance and SHAP analysis, to enhance model transparency at both the cohort level to identify key factors, and the individual level to provide personalized insights.
3 Materials and Methods
3.1 Data
This study utilized two key de-identified datasets from different healthcare organizations: administrative claims data and clinical data. The de-identified administrative claims dataset, obtained from a major health insurance provider and approved by their Institutional Review Board, includes patient interactions with healthcare professionals from January 1, 2009, to December 31, 2018. It encompasses diagnosis codes, treatment records, and healthcare costs for individuals diagnosed with Chronic Kidney Disease (CKD) and managed by a leading community nephrology practice that shared their de-identified clinical data.
The clinical dataset was sourced from the EHR of the nephrology practice, containing laboratory results, patient demographics, diagnostic details, and medication records from 1998 to 2018. For consistency of integration, the clinical data was truncated to match the 10-year span of the claims data. Usage of the clinical data was approved by the Institutional Review Board of the University.
Preprocessing of the claims data involved removing duplicates, excluding records without any stage of CKD diagnoses, and eliminating entries with negative costs to ensure data integrity. After cleansing, the claims dataset included 5,317,178 individual claims across 7,129 unique patients. The clinical dataset was similarly refined to include 433,421 lab records for 10,326 patients. These datasets were then integrated using unique, de-identified patient IDs, providing a comprehensive view of each patient’s medical and claims history. This integration was crucial for accurately identifying ESRD predictors using multiple analytical models.
3.2 Research Framework
Figure 1: Overview of research framework and training strategy. Part A provides an overview of the data preparation, modeling and validation, and additional analyses pipeline. Part B illustrates the cohort identification process, detailing the inclusion and exclusion criteria applied to claims and clinical records. Part C outlines the modeling strategy.
As illustrated in Fig.1, our framework for predicting Chronic Kidney Disease (CKD) progression to End-Stage Renal Disease (ESRD) consists of three key stages: Data Preparation, Modeling & Validation, and Additional Analyses form robustness and clinical utility.
3.2.1 Data Preparation
Data Pre-processing is the foundation of the framework, involving the integration and thorough cleaning of claims and clinical datasets. This Part A stage includes data preprocessing steps such as cleansing, normalization, missing value imputations and log transformations as appropriate for data analysis.
Cohort Identification: Following this, in Part B, cohort identification was conducted to define patient subsets aligned with the research objectives, ensuring that data included full observation windows after CKD stage 3 and excluded those who developed ESRD within this period. Although the example in the figure uses an 18-month observation window, the exact duration may vary based on the selected window of observed data.
The study cohort was curated through a multi-step process, focusing on patients diagnosed at CKD stage 3 to capture a critical period for potential intervention available to nephrologists. We retained patients with comprehensive data extending beyond the observation window, ensuring a longitudinal view of their CKD management. Further refinement excluded patients who progressed to ESRD within the observation window, resulting in a final cohort with uninterrupted, detailed medical and claims histories.
Feature Engineering, depicted in Part C, involved identifying and creating predictive variables essential for modeling. With claims data, we categorized the features used for prediction into two main groups. The first, cost-based features, was derived from five types of claims: Inpatient, Outpatient, Professional, Pharmacy, and Vision. These features include the count of unique claims per type, the aggregate cost of each type, the range of claim costs, and the standard deviation of claim costs for each patient.
The second group, comorbidity-based features, includes patient-specific clinical information such as the duration in CKD stage 3, the number of emergency department (ED) visits, and the presence of critical comorbidities such as hypertension, diabetes, and phosphatemia. These features are crucial for understanding CKD progression and were statistically analyzed for their significance in predicting ESRD outcomes (as detailed in Table 1(a) and Table 1(b)).
Table 1: Summary of Descriptive and Categorical Statistics for CKD Features
(a) Descriptive statistics for claims-driven and clinical-driven features; means and standard deviations (StD) for ESRD vs. non-ESRD populations. The $p$-value is based on the t-test of the difference between the ESRD and non-ESRD populations. A significance level of $0.05$ is used.
Category Feature Name Overall Mean (StD) ESRD Mean (StD) Non-ESRD Mean (StD) $p$-value
Claims-driven Count of pharmacy claims (CM1) 110 (88) 120 (94) 109 (87) 0.035
Count of inpatient claims (CM2) 3.74 (3.80) 3.75 (3.43) 3.74 (3.87) 0.959
Count of outpatient claims (CM3) 23.69 (22.07) 27.78 (28.75) 23.09 (20.83) 0.003
Count of professional claims (CM4) 89.73 (68.53) 105.37 (77.46) 87.43 (68.01) 0.000
Net cost of pharmacy claims (CM5) 10645 (20303) 12053 (17596) 10440 (20662) 0.110
Net cost of inpatient claims (CM6) 30038 (36080) 33909 (53540) 29440 (32541) 0.121
Net cost of outpatient claims (CM7) 8657 (17495) 9354 (17522) 8554 (17492) 0.415
Net cost of professional claims (CM8) 12137 (13707) 15512 (18657) 11640 (12748) 0.000
Range of claims costs (CM9) 9173 (15910) 11352 (32606) 8852 (11550) 0.147
Standard deviation of claims costs (CM10) 766 (879) 831 (1263) 757 (806) 0.277
Clinical-driven Age at CKD stage 3 first diagnosis (CL1) 72.68 (11) 70.13 (10.37) 73.04 (11) 0.000
CKD stage 3 duration time (days) (CL2) 660 (179) 543 (250) 677 (160) 0.000
Number of emergency department visits (CL3) 2.11 (2.94) 2.18 (2.63) 2.01 (2.99) 0.258
(b) Categorical clinical-driven features; proportions of ESRD vs. non-ESRD populations. The difference between the ESRD and non-ESRD populations is statistically significant for most of the clinical-driven categorical features based on chi-squared test of independence results with a significance level of $0.05$.
Category Feature Name ESRD Proportion Non-ESRD Proportion ${\chi}^{2}$ Statistics $p$-value
Clinical-driven Gender (CL4) (male:female) 60:40 52:48 6.73 0.009
Occurrence of CKD stage 4 (CL5) 43:57 12:88 217.8 0.000
Occurrence of CKD stage 5 (CL6) 5:95 1:99 48.8 0.000
Diabetes (CL7) 73:27 59:41 27.3 0.000
Anemia (CL8) 64:36 62:38 0.58 0.445
Metabolic acidosis (CL9) 25:75 18:82 9.41 0.002
Proteinuria (CL10) 13:87 17:83 3.72 0.054
Secondary hyperparathyroidism (CL11) 33:67 18:82 42.9 0.000
Phosphatemia (CL12) 5:95 3:97 3.55 0.06
Atherosclerosis (CL13) 6:94 14:86 18.04 0.000
Heart failure (CL14) 5:95 11:89 13.7 0.000
Stroke (CL15) 1:99 3:97 4.58 0.032
Conduction & dysrhythmias (CL16) 5:95 16:84 29.9 0.000
Hypertension (CL17) 99:1 97:3 3.1 0.078
Clinical data features were categorized into three groups based on consultations with nephrologists, commonly referenced literature, and the characteristics of our datasets. These groups include demographic features, laboratory data features, and comorbidity features.
  • Demographic Features: These include age, gender, ethnicity and BMI, which have been widely used for understanding CKD progression.
  • Laboratory Data Features: This group consists of critical laboratory results such as eGFR, hemoglobin, phosphorus, serum calcium, bicarbonate, and albumin, which are important for monitoring CKD. Notably, UACR was not selected due to insufficient readings compared to the other features.
  • Additional Comorbidity Features: These capture the presence of additional health conditions such as cardiovascular disease, anemia, and metabolic acidosis, in addition to those identified using claims data, and have been shown to influence CKD outcomes.
3.2.2 Modeling & Validation
To develop predictive models for CKD progression to ESRD, we employed both machine learning (ML) and deep learning (DL) frameworks to capture complex data patterns. Our modeling strategy (Fig. 1 Part C) involved partitioning data into training (60%), validation (20%), and testing (20%) sets to ensure model reliability. We optimized performance through hyper-parameter tuning and an iterative process to determine the optimal observation window that balances predictive accuracy and timeliness.
Machine Learning Methods
We employed statistical methods, beginning with logistic regression (LR) 11 as a baseline to estimate the probability of ESRD progression using clinical and claims data features. Beyond LR, we utilized ML techniques such as Random Forest (RF) 12 and Extreme Gradient Boosting (XGBoost) 13. RF uses ensemble learning by constructing multiple decision trees, which improves model accuracy and mitigates overfitting. XGBoost, a gradient boosting algorithm, builds models iteratively to correct errors from previous iterations, proving highly effective for structured data.
To enhance model interpretability, we conducted cohort-level feature importance analysis to identify key predictors across the dataset. Additionally, SHAP analysis was applied at the individual level to elucidate each feature’s contribution to specific predictions. This dual approach—cohort-level and individual-level analysis—not only provides insights into the drivers of ESRD progression but also improves model transparency, supporting personalized treatment decisions.
Deep Learning Methods
To develop a robust predictive model for CKD progression to ESRD, we followed methodologies from prior studies 14; 15 and constructed a longitudinal data representation segmented into three-month intervals from each patient’s initial CKD stage 3 diagnosis. This approach captures the temporal evolution of CKD, reflecting its gradual progression and aligning with clinical guidelines for periodic nephrologist visits, while also addressing potential time lags between claims and clinical data. Numerical features were aggregated, categorical features were encoded binarily to capture conditions or events, and a baseline assessment of historical records was conducted to adjust for pre-existing conditions, ensuring the model accurately accounted for CKD presence and progression.
We evaluated several deep learning architectures: Convolutional Neural Network (CNN) 16 for feature extraction, Recurrent Neural Network (RNN) 17, Long Short-Term Memory network (LSTMs 18, Gated Recurrent Unit (GRU) 19 for temporal dependency modeling, and Temporal Convolutional Network (TCN) 20 for sequence modeling. These models were chosen to provide a robust foundation for accurately predicting ESRD progression, identifying key risk factors, and anticipating CKD to ESRD transition with high precision.
Modeling Scenarios: Robustness and Clinical Application
As outlined in the Objective section, our modeling approach addresses two key scenarios:
  • Model Robustness: To ensure model consistency and fair comparability, we train and test on a unified dataset. We address the issue of shrinking cohort size with longer observation windows by selecting cohorts with a minimum of 30 months of data following CKD stage 3 diagnosis. The dataset is partitioned into training (60%), validation (20%), and testing (20%) subsets. By varying training data length from 6 to 30 months within this cohort, we assess model performance under different data availability conditions, thereby evaluating robustness and reliability.
  • Clinical Utility: In clinical practice, the progression timeline from CKD to ESRD is often uncertain. To model this uncertainty, we integrate all time-length cohorts, enabling the prediction of ESRD likelihood and optimal observation windows without precise progression timelines. This mixed-cohort approach enhances model relevance in real-world clinical settings, accommodating variable patient timelines.
3.2.3 Additional Analyses
This stage includes a detailed examination of incorrect predictions, identification of factors contributing to errors, and exploration of potential biases, particularly considering the recent changes to eGFR equations in 2021. These analyses are critical for refining the models and ensuring they deliver fair and accurate predictions.
Analysis of Model Misclassifications
To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set. This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes.
We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error). By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
This detailed error analysis is crucial for improving the model’s accuracy and ensuring its reliability in clinical settings. By understanding the reasons behind incorrect predictions, we can enhance the model’s ability to support clinicians in making more informed and accurate decisions regarding patient care.
Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias. In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare. In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 21, which excluded the race coefficient to improve accuracy and equity 22. This adjustment aimed to enhance prediction precision, especially for minority populations 23; 24; 21.
Our analysis, using data from 2009 to 2018, initially applied the older eGFR equation. To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
4 Results
4.1 Data Integration and Cohort for Analysis
(a) Comparison of CKD stage distribution across claims, clinical, and merged datasets with a 24-month observation window.
(b) Variation in CKD stage distribution as a function of different observation window lengths.
Figure 2: Overview of CKD Stage Distribution in Merged Data Across Observation Window Lengths.
To maximize the utility of multi-sourced integrated data, we merged claims and clinical datasets using unique patient identifiers. The distribution of data for individual and merged datasets is presented in Fig. 2. The analysis cohort consists of patients identified at CKD stage 3, marking the beginning of data collection for each patient. As shown in Fig. 2(a), merging the datasets resulted in a significant reduction in the number of patients—from 5,518 and 6,842 in the claims and clinical datasets, respectively, to 1,422 in the merged dataset with a 24-month observation window. This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis. Fig. 2(b) illustrates the changes in the total number of patients and their distribution across CKD stages with varying observation window lengths. As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis. This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period. Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
4.2 Performance Comparison of Models Across Datasets and Scenarios
4.2.1 Scenario 1: Model Robustness
This scenario focuses on evaluating the robustness and reliability of various models by training and testing them on a consistent dataset. We varied the training data length from 6 to 30 months within a fixed cohort to assess the models’ performance under different data availability conditions. This approach enabled us to evaluate the models’ stability and effectiveness across various observation windows, ensuring consistency and fair comparability.
Table 2: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 1 (Model Robustness). This evaluation focuses on the stability and effectiveness of models across varying observation window lengths. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
(a) Performance of models using claims data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Claims Data Only Modeling Machine Learning Methods Logistic Regression 0.61 0.24 0.64 0.31 0.70 0.32 0.72 0.33 0.71 0.29
Random Forest 0.69 0.26 0.73 0.33 0.74 0.41 0.74 0.36 0.68 0.33
XGBoost 0.70 0.27 0.75 0.35 0.78 0.42 0.75 0.39 0.69 0.38
Deep Learning Methods CNN 0.72 0.36 0.73 0.42 0.81 0.44 0.82 0.45 0.84 0.44
RNN 0.77 0.39 0.75 0.44 0.83 0.49 0.90 0.50 0.82 0.47
LSTM 0.76 0.33 0.79 0.43 0.86 0.45 0.92 0.54 0.87 0.44
GRU 0.70 0.34 0.78 0.46 0.88 0.45 0.91 0.50 0.86 0.42
TCN 0.72 0.36 0.81 0.37 0.85 0.46 0.88 0.52 0.85 0.43
(b) Performance of models using clinical data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Clinical Data Only Modeling Machine Learning Methods Logistic Regression 0.63 0.46 0.70 0.49 0.76 0.51 0.76 0.54 0.75 0.53
Random Forest 0.70 0.53 0.76 0.53 0.80 0.58 0.79 0.58 0.77 0.54
XGBoost 0.73 0.52 0.76 0.55 0.80 0.59 0.80 0.57 0.72 0.55
Deep Learning Methods CNN 0.73 0.51 0.75 0.54 0.80 0.55 0.84 0.56 0.80 0.52
RNN 0.74 0.53 0.77 0.57 0.82 0.60 0.85 0.61 0.79 0.55
LSTM 0.77 0.54 0.78 0.56 0.84 0.63 0.88 0.60 0.82 0.57
GRU 0.80 0.54 0.77 0.57 0.83 0.61 0.87 0.60 0.84 0.53
TCN 0.72 0.50 0.73 0.53 0.80 0.57 0.83 0.61 0.81 0.55
(c) Performance of models using merged data.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Merged Data Modeling Machine Learning Methods Logistic Regression 0.65 0.46 0.69 0.47 0.76 0.54 0.75 0.55 0.78 0.54
Random Forest 0.74 0.44 0.77 0.52 0.83 0.58 0.84 0.60 0.80 0.55
XGBoost 0.73 0.45 0.76 0.50 0.82 0.58 0.85 0.61 0.82 0.59
Deep Learning Methods CNN 0.70 0.45 0.76 0.46 0.77 0.55 0.80 0.56 0.82 0.55
RNN 0.75 0.51 0.80 0.58 0.86 0.60 0.87 0.62 0.85 0.61
LSTM 0.76 0.53 0.84 0.59 0.93 0.63 0.93 0.65 0.84 0.59
GRU 0.75 0.51 0.76 0.56 0.91 0.62 0.90 0.63 0.86 0.57
TCN 0.71 0.49 0.74 0.57 0.88 0.60 0.89 0.61 0.91 0.59
Table 2(c) compares model performance across claims data-only scenario, clinical data-only, and merged data. In claims data-only, ML model XGBoost achieved the highest AUC (0.78) at 18 months, while Random Forest had the best F1 score (0.41). DL based LSTM outperformed all others with an AUC of 0.92 and an F1 score of 0.54 at 24 months. In clinical data-only scenario, Random Forest led with an AUC of 0.80 and an F1 score of 0.58, but LSTM again excelled with an AUC of 0.88 and an F1 score of 0.63. Merged data further improved results, with XGBoost reaching an AUC of 0.85, and LSTM and GRU achieving AUCs of 0.93 and 0.91, with F1 scores of 0.65 and 0.63. Overall, merged data and deep learning methods, particularly LSTM and GRU, delivered the best performance, with the best observation window being 24 months.
4.2.2 Scenario 2: Clinical Application
In this scenario, we addressed the challenge of predicting ESRD development in real-world clinical settings, where the duration of CKD progression is often uncertain. By examining all cohort observation windows, the models predict ESRD likelihood and determine the optimal observation window despite unknown progression timelines. This mixed cohort approach enhances the model’s relevance for clinical practice, accommodating variable patient timelines.
Table 3(c) shows model performance across claims-only, clinical-only, and merged data. In claims-only, XGBoost had the highest AUC (0.76) at 12 months among ML models, while Random Forest had the best F1 score (0.40) at 18 months. However, LSTM outperformed others with an AUC of 0.90 and an F1 score of 0.51 at 24 months. For clinical-only, Random Forest led ML models (AUC 0.79, F1 0.56 at 18 months), but LSTM was superior (AUC 0.85, F1 0.61 at 24 months). Merged data yielded the best results, with XGBoost leading ML methods (AUC 0.84 at 18 months), and LSTM and GRU excelling overall (LSTM: AUC 0.90, F1 0.60; GRU: F1 0.63 at 24 months). Merged data consistently outperformed individual datasets, and deep learning models, especially LSTM and GRU, showed significant performance gains, highlighting the benefits of integrating multiple data sources for accurate clinical predictions.
Comparing the two scenarios, we observe that while both benefit from deep learning models and data integration, they demonstrate different strengths depending on the context. Scenario 1 underscores the importance of stability and consistency across varying data conditions, whereas Scenario 2 highlights adaptability and practical applicability in real-world clinical settings. Together, these scenarios underscore the critical role of deep learning and comprehensive data integration, with 24 months identified as a key observation window that balances robustness and clinical relevance.
Table 3: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 2 (Clinical Application). The model aims to predict the likelihood of ESRD development and suggest an optimal observation window based on data from the patient’s initial stage 3 CKD visit. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).
(a) Performance of models using claims data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Claims Data Only Modeling Machine Learning Methods Logistic Regression 0.66 0.22 0.67 0.28 0.70 0.32 0.72 0.34 0.71 0.35
Random Forest 0.72 0.26 0.75 0.30 0.74 0.40 0.73 0.34 0.68 0.31
XGBoost 0.72 0.28 0.76 0.31 0.72 0.36 0.71 0.32 0.69 0.33
Deep Learning Methods CNN 0.70 0.34 0.72 0.48 0.79 0.49 0.88 0.46 0.86 0.42
RNN 0.79 0.43 0.78 0.47 0.82 0.47 0.84 0.49 0.82 0.41
LSTM 0.79 0.33 0.88 0.33 0.84 0.47 0.90 0.51 0.87 0.41
GRU 0.73 0.29 0.85 0.44 0.86 0.50 0.89 0.49 0.86 0.40
TCN 0.72 0.42 0.82 0.38 0.84 0.48 0.87 0.49 0.85 0.40
(b) Performance of models using clinical data only.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Clinical Data Only Modeling Machine Learning Methods Logistic Regression 0.61 0.45 0.69 0.48 0.78 0.49 0.76 0.50 0.73 0.51
Random Forest 0.72 0.51 0.75 0.53 0.79 0.56 0.73 0.54 0.68 0.50
XGBoost 0.72 0.51 0.76 0.54 0.78 0.58 0.71 0.56 0.69 0.57
Deep Learning Methods CNN 0.70 0.44 0.72 0.46 0.77 0.46 0.80 0.56 0.76 0.54
RNN 0.74 0.48 0.72 0.44 0.78 0.48 0.82 0.59 0.78 0.54
LSTM 0.76 0.44 0.78 0.46 0.80 0.51 0.85 0.61 0.79 0.55
GRU 0.73 0.46 0.78 0.52 0.80 0.50 0.85 0.59 0.77 0.47
TCN 0.69 0.47 0.75 0.48 0.77 0.48 0.81 0.59 0.77 0.51
(c) Performance of models using merged data.
6 months 12 months 18 months 24 months 30 months
Model/Metrics AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
Merged Data Modeling Machine Learning Methods Logistic Regression 0.66 0.44 0.68 0.47 0.74 0.51 0.72 0.51 0.65 0.48
Random Forest 0.72 0.47 0.76 0.51 0.80 0.56 0.81 0.52 0.77 0.53
XGBoost 0.71 0.50 0.79 0.52 0.84 0.57 0.81 0.56 0.79 0.57
Deep Learning Methods CNN 0.60 0.43 0.66 0.43 0.77 0.47 0.78 0.48 0.76 0.50
RNN 0.66 0.49 0.69 0.44 0.80 0.50 0.84 0.55 0.80 0.52
LSTM 0.68 0.49 0.75 0.48 0.89 0.57 0.90 0.60 0.79 0.53
GRU 0.65 0.49 0.76 0.49 0.89 0.55 0.90 0.63 0.79 0.52
TCN 0.65 0.45 0.74 0.50 0.78 0.44 0.88 0.59 0.78 0.53
4.3 Key Factor Analysis Using Explainable AI Techniques
We identified a 24-month observation window as optimal for balancing early detection and prediction accuracy, with LSTM proving most effective among deep learning models and XGBoost leading in machine learning. Focusing on Scenario 2 for its relevance to real-world clinical settings with uncertain CKD progression timelines, we refined our approach further to enhance clinical applicability. We used cohort-level feature importance analysis with XGBoost to determine key predictive features and applied SHAP analysis at the individual patient level to improve model interpretability and transparency.
4.3.1 Cohort-Level Key Feature Identification
Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.  3(a)). The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its significance in predicting ESRD. Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly. The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power. This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice. However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
(a) Top 30 features of importance for the XGBoost model.
(b) SHAP analysis demonstrating the variation in feature impact on ESRD risk prediction for two patients with CKD Stage 5.
Figure 3: Key Feature Analysis: Feature Importance and SHAP Analysis for ESRD Risk Prediction Using A 24-month Observation Window.
4.3.2 Feature Impact at the Individual Patient Level Using SHAP Analysis
While the previous subsection explored feature importance at the cohort level, the uniqueness of each patient’s condition necessitates a more personalized approach. To aid healthcare providers in making tailored decisions, we applied SHAP analysis to identify the key features driving individual patient predictions.
We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD. Despite both having CKD Stage 5, the contributing features differed significantly, underscoring the need for personalized treatment strategies.
In Figure 3(b), features increasing ESRD risk are highlighted in red, while those reducing risk are in blue. For Patient X, factors such as Stage 5 CKD (S5) and a high number of outpatient claims (n_claims_DR_min) elevated the risk, while higher eGFR levels and the absence of certain markers reduced it. For Patient Y, lower risk was driven by factors such as younger age, fewer minimum inpatient claims (net_exp_O_min), and lower minimum eGFR values, despite also having a Stage 5 CKD diagnosis (S5).
This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning. SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile, thereby enhancing the effectiveness of clinical care.
4.4 Analysis of Model Misclassifications: Type I and Type II Errors
To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 4(a). Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered. This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
To understand what drives these misclassifications, we conducted a detailed analysis of Type I and Type II errors, illustrated in Figures 4(c) and 4(b). Type II errors (Figure 4(b)) involve patients wrongly predicted not to develop ESRD but did; we compared these 16 false negatives to the 70 correctly predicted to progress to ESRD. These errors can delay essential treatment, potentially accelerating disease progression. Conversely, Type I errors (Figure 4(c)) involve patients wrongly predicted to develop ESRD who did not. We contrasted 17 false positives with 1,319 correct predictions of non-progression, highlighting how these errors can lead to unnecessary medical interventions, potentially depriving resources from those who truly need them. By examining the feature distributions between these correctly and incorrectly predicted groups, we aimed to identify statistically significant differences that could explain the model’s errors, providing insights for refining the model and improving its predictive accuracy.
(a) Distribution of Predicted Probabilities for Patient Outcome.
(b) Analysis of Type II Errors: Comparison of Patients Predicted Not to Develop ESRD Who Actually Did.
(c) Analysis of Type I Errors: Comparison of Patients Predicted to Develop ESRD Who Actually Did Not.
Figure 4: Analysis of Model Misclassifications: Type I and Type II Errors. The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset. The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
Table 4: Analysis of Type I and Type II Errors: Features and Their Impact on Prediction Accuracy. This table presents representative subsets of features under various timestamps. Features in bold appear in both analyses.
(a) Type II Error Analysis: Feature Comparison Between False Negatives and True Positives.
Model Mean P-value Significant Timestamp
Correct incorrect
n_claims_I 0.70 0.14 0.048752 Yes 5
n_claims_O 5.71 2.67 0.001325 Yes 2
net_exp_O 3831.83 830.61 0.002747 Yes 0
S5 0.36 0.06 0.001189 Yes 6
S5 0.37 0.49 0.662073 No 7
net_exp_O 4219.01 1044.37 0.001883 Yes 4
(b) Type I Error Analysis: Feature Comparison Between False Positives and True Negatives.
Model Mean P-value Significant Timestamp
Correct incorrect
n_claims_DR 13.52 7.21 0.001156 Yes 3
n_claims_O 3.03 1.69 0.009384 Yes 4
S4 3831.83 830.61 0.002747 Yes 4
S5 0.09 0.35 4.059934e-02 Yes 6
S5 0.08 0.35 3.911497e-02 Yes 7
net_exp_O 1044.82 333.61 0.000324 Yes 4
Based on Table 4(a), CKD Stage 5 emerges as a critical feature. For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, significantly lower than the 0.36 mean for those correctly predicted to progress. This suggests that these patients did not have a Stage 5 record until the subsequent timestamp. The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely. However, after consulting with nephrologists and examining the literature on CKD progression to ESRD, we understand that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
Based on Table 4(b), CKD Stage 5 appears in both timestamps 6 and 7 for the incorrectly predicted group, indicating that these patients had Stage 5 records at these timestamps but did not actually progress to ESRD. This is unusual. One possible explanation could be the issue of data censoring. Our dataset spans from 2009 to 2018, so patients incorrectly predicted to develop ESRD, with continuous Stage 5 records, might have progressed to ESRD in 2019 or later. Additional data collection beyond 2018 would be needed to update and refine this analysis.
4.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
(a) Incorrect predictive probabilities using the 2009 eGFR equation.
(b) Incorrect predictive probabilities using the 2021 eGFR equation.
(c) Distribution of race, gender, and ESRD status with the 2021 eGFR equation. Categories: Race (0 = White, 1 = Black, 2 = Other), Gender (0 = Male, 1 = Female), ESRD (0 = No ESRD, 1 = ESRD).
(d) Detailed racial distribution of patients incorrectly predicted not to develop ESRD, comparing 2009 and 2021 eGFR equations.
Figure 5: Impact of Updated eGFR Equations (2009 vs. 2021) on Prediction Errors and Racial Bias. Subplots (a) and (b) compare incorrect predictions using the two eGFR equations, while (c) and (d) explore the demographic distribution and racial disparities among incorrectly predicted patients.
When comparing the incorrect predictions using the 2009 eGFR equation (Fig  5(a)) with those from the 2021 equation (Fig  5(b)), we observed a decrease in the number of incorrect predictions from 33 to 28. This reduction reflects a slight improvement in the error rate, which decreased from 0.0232 to 0.0228 with the updated 2021 eGFR equation. Despite this improvement, the pattern of incorrect predictions being clustered towards the lower end, as discussed earlier, persists.
To assess whether the updated eGFR equation improves prediction accuracy for minoritized groups, we analyzed the demographic and racial distribution of patients who were incorrectly predicted not to develop ESRD. Fig  5(c) displays the demographic breakdown, including race, gender, and ESRD outcomes, for the newly defined cohort under the 2021 equation. Fig  5(d) reveals changes in incorrect predictions: the number of misclassified white patients increased from 2 to 4, while incorrect predictions for African American patients decreased significantly from 4 to 1. Meanwhile, the count for patients in the ”Other” category rose from 10 to 12.
Given the relatively small sample size for the ”Other” racial category, the increase may not be statistically significant. However, the notable reduction in incorrect predictions for African American patients suggests that the 2021 eGFR equation enhances the model’s prediction accuracy for this group. This improvement is particularly important as it indicates progress towards reducing racial bias in ESRD predictions, aligning with broader efforts to enhance equity in healthcare.
Overall, while the updated 2021 eGFR equation has led to a modest reduction in the overall error rate and appears to improve prediction accuracy for African American patients, further analysis with larger cohorts are needed to assess its impact on other racial groups. This finding underscores the ongoing need to refine prediction models to ensure fair and accurate outcomes across all patient demographics.
5 Discussion
This study underscores the value of integrating patient data from multiple sources, such as claims and EHRs, and of different types, such as clinical comorbidities and utilization costs, combined with deep learning methodologies, particularly in the longitudinal representation of healthcare data, for improving the accuracy of patient outcome predictions. By employing models such as LSTM and GRU, which are adept at capturing temporal dependencies in sequential data, we were able to enhance the predictive power to assess progression to ESRD in CKD patients. The use of a 24-month observation window, identified as the optimal period for accurate predictions, allows for timely and reliable risk assessment, providing a critical balance between early detection and maintaining a manageable error rate. These findings add to the rapidly growing evidence for driving more precise and interpretable predictions for chronic disease management, ultimately supporting better clinical decision-making to improve patient outcomes.
This research bridges a significant gap in risk assessment for chronic disease management studies by integrating clinical and claims data. Previous models have often relied on one data type in isolation, either clinical or claims, which can lead to incomplete representations of a patient’s health journey and the progression of CKD. By combining both data types, our approach offers a more comprehensive view, capturing not only medical indicators but also healthcare utilization patterns, as demonstrated by the features that are identified as most important in ESRD prediction. This integrated data strategy enhances the model’s robustness and applicability across diverse patient populations, addressing socio-economic factors and healthcare interactions that single-source models may overlook. Our findings highlight the importance of a multi-faceted approach to data integration, which can lead to more inclusive and accurate models for managing chronic diseases such as CKD.
A key innovation of our work is the determination of an optimal observation window that minimizes the volume of data and time required for making reliable predictions while maintaining an acceptable error rate. This focus on observation period optimization is crucial for balancing the need for timely interventions with the risks of premature or inaccurate predictions. By rigorously evaluating different observation window lengths, we were able to identify a 24-month period as optimal, providing a critical framework for clinicians to monitor and manage CKD progression effectively. This approach supports a proactive stance in patient care, allowing for earlier detection and management of patients at risk of progressing to ESRD, thereby potentially reducing long-term healthcare costs and improving patient outcomes.
The insights derived from our research have significant practical implications for early intervention strategies in CKD management. By leveraging integrated data and advanced modeling techniques, healthcare providers can better identify patients at high risk of advancing to ESRD, enabling more targeted and timely interventions. This approach aligns with the goals of personalized medicine, where treatment plans are tailored to the specific needs and risk profiles of individual patients. Moreover, the reduced error rates with the updated 2021 eGFR equation suggest an improvement in prediction accuracy for minoritized populations, particularly African American patients, which is a positive step toward reducing healthcare disparities. These findings advocate for continued efforts to refine and optimize predictive models, ensuring equitable and effective care for all patients, regardless of their demographic background.
5.1 Limitations
Our study’s reliance on data from a single institution limits the model’s generalizability to diverse patient populations. Incorporating data from more sources could enhance robustness and applicability across different clinical settings. The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
The use of Electronic Health Records (EHR) data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness. Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
Additionally, the model does not fully account for time lags between claims and clinical data, which may affect the precision of predictions. Reducing these discrepancies is essential for enhancing the model’s effectiveness in clinical practice.
5.2 Future Directions
Addressing Data Censoring: To mitigate data censoring issues, an initial strategy could involve truncating the data up to 2016 and analyzing patient outcomes in 2017 and 2018. Following this, expanding the dataset to include more recent patient information beyond 2018 would be crucial. This expansion would enhance our understanding of health trajectories and disease progression, enabling continuous refinement of the model to better align with evolving healthcare practices and patient demographics.
Integration of Unstructured Clinical Notes: Future model iterations will integrate unstructured clinical notes, such as physician narratives and discharge summaries, to capture detailed patient information often missed in structured formats. This approach aims to provide a more comprehensive view of patient health, enhancing outcome predictions and enabling more personalized interventions.
Advanced Methods for Time-Lag Mitigation: Advanced techniques to mitigate time lags between claims and clinical data, such as algorithms that adjust for expected delays or incorporate more real-time data sources for more accurate triangulation, are needed to improve the model’s ability to handle these temporal discrepancies, enhance prediction accuracy and ensure timely, data-driven interventions.
Expansion to Other Chronic Diseases: To evaluate our framework’s versatility, we plan to apply it to other chronic conditions, such as diabetes and heart disease. This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
6 Conclusion
As multiple sources of patient-level healthcare data continue to get integrated and analyzed in large data repositories and cloud computing platforms, this study presents an insightful illustrative approach for predicting ESRD outcome in CKD patients by integrating clinical and claims data with advanced machine learning and deep learning methods. Our results demonstrate that this combined data strategy significantly improves prediction accuracy, offering a more holistic understanding of CKD progression. Through explainable AI techniques such as SHAP analysis, we identified critical factors influencing predictions at both individual and cohort levels.
A key contribution of this work is the development of a framework that optimizes the observation window, balancing early detection with predictive accuracy to enhance clinical decision-making. The 24-month observation window was identified as the optimal period for maximizing prediction effectiveness while minimizing unnecessary interventions. Additionally, our assessment of the updated 2021 eGFR equation highlights an approach to health equity, contributing to fairer healthcare outcomes.
Overall, this study advances chronic disease management by utilizing integrated data and innovative AI techniques, paving the way for more accurate, personalized, and equitable care in CKD management. These findings lay the groundwork for future research to refine predictive models and expand their application to other chronic diseases.
7 Acknowledgments
We sincerely thank the community nephrology practice for providing access to their de-identified clinical data and the wealth of medical knowledge and inputs shared by several nephrologists in the practice and the CIO on the raw data and the results of our analysis. Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice, partial funding for the study, and the invaluable help provided by their data team in understanding the claims data. We also acknowledge the fellowship support for Y.L from the Center for Machine Learning and Health (CMLH) at Carnegie Mellon University.
8 Conflicts of interest
None declared.